find 3 consecutive even numbers such as that sum of the first and the third consecutive even numbers is equal to the second increased by 50
Found 2 solutions by Explosionzz, mananth:
Answer by Explosionzz(1)

x=1st integer
x+1=2nd integer
x+2=3rd integer
x+(x+2)=(x+2)+50
2x+2=x+52
x+2=52
x=50
50, 51, and 52.

Answer by mananth(16949)

find 3 consecutive even numbers such as that sum of the first and the third consecutive even numbers is equal to the second increased by 50
...
let the consequtive even numbers be x,x+2,x+4
x+x+4=x+2+50
2x+4=x+52
2x-x=48
x=48
..
the numbers are 48,50,52
...
m.ananth@hotmail.ca